OpenStudy (anonymous):

A line has a slope of 5/6 and a y-intercept of –3. Which is the equation of the line?

OpenStudy (anonymous):

Option A: y = -3x + 5/6 Option B: y = 3x + 5/6 Option C: y = 5/6x-3 Option D: y=5/6x+3

OpenStudy (lgbasallote):

slope intercept is y = mx + b m is slope b is y intercept does that help?

OpenStudy (anonymous):

is 5/6 the slope?

OpenStudy (lgbasallote):

yes

OpenStudy (lgbasallote):

and y intercept is?

OpenStudy (anonymous):

-3

OpenStudy (lgbasallote):

right. so slope intercept form is?

OpenStudy (anonymous):

option D?

OpenStudy (anonymous):

@igbasallote right or wrong..

OpenStudy (lgbasallote):

wrong. y intercept is -3 so b should be -3